OpenStudy (anonymous):

Partial Fractions\[\int\limits_{}^{}\frac{ dx }{ x^2(x^2-16) }\]

OpenStudy (anonymous):

should I split up x^2-16

hartnn (hartnn):

without partial fraction, there's a way, adjust the numerator, 1 = -16/-16 1 = (-1/16) (x^2-16 -x^2) did you get this adjustment ?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

i get the top part but not after that

hartnn (hartnn):

we want numerator = x^2-16 -x^2 {=-16 } so that we can split the denominator but the actual numerator =1 so i just multiplied and divided by -16

OpenStudy (anonymous):

if you want to do partial fractions, then yes, you must split up (x^2 - 16) into (x+4)(x-4)

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

but what hartnn is doing is pretty cool too. \[\int\limits_{}^{} \frac{ dx }{ x^2(x^2 - 6) } \int\limits_{}^{}\left( -\frac{ 1 }{ 16 } \right)\frac{ x^2 - 16 - x^2 }{ x^2(x^2 - 16) } dx = -\frac{ 1 }{ 16 }\int\limits_{}^{}\frac{ 1 }{ x^2 } - \frac{ 1 }{ x^2 - 16 } dx\]

OpenStudy (anonymous):

forgot the first equal sign lol

OpenStudy (anonymous):

with partial fractions: \[\int\limits_{}^{}\frac{ dx }{ x^2(x^2 - 16) } =>\frac{ 1 }{ x^2(x-4)(x+4) }= \frac{ A }{ x } + \frac{ B }{ x^2 } + \frac{ C }{ x-4 } + \frac{ D }{ x+4 }\] so much longer than adjusting the numerator.

OpenStudy (anonymous):

ok thanks you guys